OpenStudy (anonymous):

Help!!! me!!! A line has slope 8 and (3, -1) is on the line. Which is the correct equation that describes this line in slope-intercept form? A:8x + 25 = y B:8x + y = 25 C:y = 8x – 25 D:y + 25 = 8x

OpenStudy (anonymous):

y=mx+b m=8 x=3 y=-1 -1=8(3)+b b= -25 therefore : y=8x-25

OpenStudy (anonymous):

Thank you!